OpenStudy (anonymous):

if I were to get $100.00 of a $250.00 prize, what percent of that prize would I be getting?

OpenStudy (anonymous):

40%

OpenStudy (anonymous):

100/250 * 100 gives percentage , solve it

OpenStudy (anonymous):

= 10000/250 = 40 percent

OpenStudy (anonymous):

his is getting 100 out of 250 so (100/250) what percent (per -cent ie per 100) so (100/250) x 100 answer will be 40 %

OpenStudy (anonymous):

if half (50% is 125.00 ) what is the answer?

OpenStudy (anonymous):

let x be the amount 50% of x = 125 50/100 * x = 125 x = 125 * 100/50 = 250

OpenStudy (anonymous):

Prize earned / Prize total = Decimal Decimal times 100 = Percent 100/250 = 0.40 0.40 * 100 = 40% Answer is 40%